# Regularized KL-Divergence for Well-Defined Function-Space Variational Inference in Bayesian Neural Networks

### v2

@@ A.3.1 Experiments on synthetic data
$y_{i}=\operatorname{sin}(2\pi x_{i})+\epsilon\quad\text{with}\quad\epsilon\sim\mathcal{N}\mathopen{}\left(0,\sigma_{n}^{2}\right)\mathclose{}$ (A.6)
and draw $x_{i}\sim\mathcal{U}([-1,-0.5]\cup[0.5,1])$. When not otherwise specified, we use $\sigma_{n}=0.1$. On the plots, the data points are shown as gray circles, inferred mean functions as red lines, their 2-standard-deviations interval around the mean in light green, and functions sampled from the approximate posterior as green lines. In general, we consider two hidden-layer BNNs with 30 neurons per layer and hyperbolic tangent activation (Tanh) functions. Specifically in figure 11, the small BNN has the same architecture as above while the large BNN has 100 neurons per layer. All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior. The prior scale of TFSVI (Rudner et al., 2022b, ) is set to $\sigma_{p}=0.2$ and $\sigma_{p}=0.75$ for MFVI (Blundell et al.,, 2015) and Laplace (Immer et al.,, 2021). For the Gaussian process posterior baseline, we fit the prior parameters by maximizing the log-marginal likelihood (Williams and Rasmussen,, 2006). Apart from the cases where the parameters of the GP prior used for GFSVI (our method) and FVI (Sun et al.,, 2019) are explicitly stated, we consider a constant zero-mean function and find the parameters of the covariance function by maximizing the log-marginal likelihood from mini-batches (Chen et al.,, 2021). Except where otherwise stated, we estimate the functional KL divergences with 500 measurement points and use the regularized KL divergence with $\gamma=10^{-10}$.
@@
We sample 100 data points perturbed by Gaussian noise with $\sigma_{n}=0.1$ from the two moons data (Pedregosa et al.,, 2011). On the plots, the data points are shown as red (class 0) and blue (class 1) dots. We plot the mean and 2-standard-deviations of the probability that ${\bm{x}}$ belongs to class 1 with respect to the posterior (i.e. $p(y=1\,|\,{\bm{w}}^{(k)},{\bm{x}})$) which we estimate from $K=100$ samples ${\bm{w}}^{(k)}\sim q_{\phi}({\bm{w}})$ for $k=1,\dots,K$. We consider two hidden-layer BNNs with 100 neurons per layer and hyperbolic tangent activation (Tanh) functions. All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior. The prior scale of MFVI (Blundell et al.,, 2015) is set to $\sigma_{p}=0.8$ and $\sigma_{p}=1.0$ for TFSVI (Rudner et al., 2022b, ) and Laplace (Immer et al.,, 2021). For the Gaussian process posterior baseline, we approximate the intractable posterior using the Laplace approximation and find the prior parameters by maximizing the log-marginal likelihood (Williams and Rasmussen,, 2006). The GP prior for GFSVI (our method) and FVI (Sun et al.,, 2019) has a constant zero-mean function and we find the parameters of the covariance function by maximizing the log-marginal likelihood from mini-batches (Chen et al.,, 2021) using the method to transform classifications labels into regression targets from Milios et al., (2018). We estimate the functional KL divergences with 500 measurement points and use the regularized KL divergence with $\gamma=10^{-10}$.
A.3.2 Ocean current modeling experiment
Following Cinquin et al., (2024), we apply the Helmholtz decomposition to the neural network $f$ as
$f(\cdot,{\bm{w}})=\operatorname{grad}\Phi(\cdot,{\bm{w}}_{1})+\operatorname{rot}\Psi(\cdot,{\bm{w}}_{2})$ (A.7)
@@ A.3.2 Ocean current modeling experiment
A.3.3 Regression experiments with tabular data
@@
We evaluate the predictive performance of our model on regression datasets from the UCI repository (Dua and Graff,, 2017) described in table 6. These datasets are also considered in Sun et al., (2019); Wild et al., (2022) but we include two additional larger ones (Wave and Denmark). We perform 5-fold cross validation, leave out one fold for testing, consider 10% of the remaining 4 folds as validation data and the rest as training data. We report mean and standard-deviation of the average expected log-likelihood and average mean square error on the test fold. We also report the mean rank of the methods across all datasets by assigning rank 1 to the best scoring method as well as any method who’s error bars overlap with the highest score’s error bars, and recursively apply this procedure to the methods not having yet been assigned a rank. The expected log-likelihood is estimated by Monte Carlo integration when it is not available in closed form (MFVI, TFSVI and FVI) with 100 posterior samples. We preprocess the dataset by encoding categorical features as one-hot vectors and standardizing the features and labels.
@@
We compare our GFSVI method to two weight-space inference methods (mean-field variational inference (Blundell et al.,, 2015) and linearized Laplace (Immer et al.,, 2021)) and two function-space inference methods (FVI (Sun et al.,, 2019) and TFSVI (Rudner et al., 2022b, )). While FVI uses GP priors, TFSVI performs inference in function space but with the pushforward to function space of the variational distribution and prior on the weights. We compute the function-space (regularized) KL divergence using a set of 500 measurement points sampled from a uniform distribution for GFSVI and TFSVI, and 50 points drawn from a uniform distribution along with 450 samples from the training batch for FVI as specified in Sun et al., (2019). All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior. We also provide results with a GP (Williams and Rasmussen,, 2006) when the size of the dataset allows it, and a sparse GP (Hensman et al.,, 2013). As we restrict our comparison to BNNs, we do not consider the GP and sparse GP as baselines but rather as gold-standards. All models have a Gaussian heteroskedastic noise model with a learned scale parameter. All the BNNs are fit using the Adam optimizer (Kingma and Ba,, 2017) using a mini-batch size of 2000 samples. We also perform early stopping when the validation loss stops decreasing.
@@
Hyper-parameter optimization is conducted using the Bayesian optimization tool provided by Wandb (Biewald,, 2020). BNN parameters are selected to maximize the average validation expected log-likelihood across the 5 cross-validation folds. We optimize over prior parameters (kernel and prior scale), learning-rate and activation function. We select priors for GFSVI, FVI, sparse GP and GP among the RBF, Matern-1/2, Matern-3/2, Matern-5/2, Linear and Rational Quadratic covariance functions. The GP prior parameters used with GFSVI and FVI are selected by maximizing the log-marginal likelihood from batches as proposed by Chen et al., (2021) and done in Sun et al., (2019). Hyper-parameters for GPs and sparse GPs (kernel parameters and learning-rate) are selected to maximize the mean log-marginal likelihood of the validation data across the 5 cross-validation folds.
@@
A.3.4 Classification experiments with image data
@@
We further evaluate the predictive performance of our model on classification tasks with the MNIST (LeCun et al.,, 2010) and Fashion MNIST (Xiao et al.,, 2017) image data sets. We fit the models on a random subset of 90% of the provided training split, consider the remaining 10% as validation data and evaluate on the provided test split. We repeat this procedure 5 times with different random seeds and report the mean and standard-deviation of the average expected log-likelihood, accuracy and expected calibration error (ECE) of the mean of the predictive distribution on the test set. The expected log-likelihood is estimated by Monte Carlo integration with 100 posterior samples when it is not available in closed form (MFVI, TFSVI and FVI). We estimate the mean of the predictive distribution to compute the accuracy and the ECE with 100 posterior samples. We preprocess the dataset by standardizing the images.
@@
We compare our GFSVI method to the same baselines as for the regression experiments (see A.3.3). All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior. More specifically, we consider a CNN with three convolutional layers (with output channels 16, 32 and 64) before two fully connected layers (with output size 128 and 10). The convolutional layers use $3\times 3$ shaped kernels. Each pair of convolutional layers is interleaved with a max-pooling layer. We consider three different measurement point distributions $\rho$ to estimate the (regularized) KL divergence in GFSVI, FVI and TFSVI: random, random pixel and kmnist. The random measurement point distribution is sampled from by drawing 50% of the samples from the training data batch and 50% of the samples from a uniform distribution over $[p_{min},p_{max}]^{H\times W\times C}$, where $H$, $W$ and $C$ are respectively the height, width and number of channels of the images, and $p_{min}=v_{min}-0.5\times\Delta$ and $p_{max}=v_{max}+0.5\times\Delta$ where $\Delta=v_{max}-v_{min}$ is the difference between the minimal ($v_{min}$) and maximal ($v_{max}$) pixel values of the data set. The random pixel measurement point distribution is taken from Rudner et al., 2022b and is sampled from by randomly choosing each pixel value among the ones available from the training data batch at the same position in the $28\times 28$ pixel grid. Finally, the kmnist measurement point distribution is also taken from Rudner et al., 2022b and is drawn from by randomly sampling data points from the Kuzushiji-MNIST (KMNIST) dataset (Clanuwat et al.,, 2018). The KMNIST dataset is a collection of 70’000 gray-scale images of size $28\times 28$ which we preprocess by standardizing the images. We sample 10 measurement points when using random, 25 measurement points when using random pixel and 20 when using kmnist. All the BNNs are trained using the Adam optimizer (Kingma and Ba,, 2017) using a mini-batch size of 100. We also perform early stopping when the validation loss stops decreasing.
@@
Hyper-parameter optimization is conducted just like for the regression tasks (see A.3.3). The Gaussian process prior parameters used with GFSVI and FVI are selected by maximizing the log-marginal likelihood from batches (Chen et al.,, 2021) using the method to transform classifications labels into regression targets from Milios et al., (2018). We optimize the same hyper-parameters as for the regression experiments with the exception of the additional $\alpha_{\epsilon}$ parameter introduced by Milios et al., (2018) for the function-space VI methods with GP priors (FVI and GFSVI).
A.3.5 OOD detection
@@
Following the setup from Malinin et al., (2020) we take epistemic uncertainty to be the variance of the mean prediction with respect to samples from the posterior. We consider the test data to be in-distribution (ID) data and a subset of the song dataset (Bertin-Mahieux et al.,, 2011) of equal length and with an equal number of features as out-of-distribution (OOD) data. We use the same preprocessing as for regression as well as the same baselines with the same hyper-parameters (see section A.3.3). We first fit a model, then evaluate the extend by which the epistemic uncertainty under the model is predictive of the ID and OOD data using a single threshold obtained by a depth-1 decision tree fit to minimize the classification loss. We report the mean and standard error of the accuracy of the threshold to classify OOD from ID data based on epistemic uncertainty across the 5 folds of cross-validation. We also provide results obtained using a GP and sparse GP as gold standard.
@@
Following the setup by Osawa et al., (2019), we take the epistemic uncertainty to be the entropy of the mean of the predictive distribution with respect to samples from the posterior. We evaluate models trained on MNIST using MNIST’s test split as ID data and a subset of the training set of Fashion MNIST as OOD data. Likewise, we evaluate models trained on Fashion MNIST using Fashion MNIST’s test split as ID data and a subset of the training set of MNIST as OOD data. We use the same preprocessing as for classification, as well as the same baselines with the same hyper-parameters (see section A.3.4). We first fit a model, then evaluate the extend by which the epistemic uncertainty under the model is predictive of the ID and OOD data using a single threshold obtained by a depth-1 decision tree fit to minimize the classification loss. We estimate mean of the predictive distribution by Monte-Carlo integration using 100 posterior samples. We report the mean and standard error of the accuracy of the threshold to classify OOD from ID data based on epistemic uncertainty for the 5 models trained on different random seeds (see section A.3.4).
A.3.6 Variational measure evaluation
We evaluate our inference method by comparing the samples drawn from the exact posterior over functions with the approximate posterior obtained with our method (GFSVI). We follow the setup by Wilson et al., (2022) and we compute the average Wasserstein-2 metric between 1000 samples drawn from a GP posterior with a RBF kernel evaluated at the test points, and samples from the approximate posterior of GFSVI, sparse GP and FVI evaluated at the same points and with the same prior. We consider the Boston, Concrete, Energy, Wine and Yacht datasets for which the exact GP posterior can be computed and use the same preprocessing as for regression (see section A.3.3). We report the mean and standard error of the average Wasserstein-2 metric across the 5 folds of cross-validation. The Wasserstein-2 metric is computed using the Python Optimal Transport library (Flamary et al.,, 2021).
@@ A.3.6 Variational measure evaluation
A.3.7 Software
We use the JAX (Bradbury et al.,, 2018) and DM-Haiku (Hennigan et al.,, 2020) Python libraries to implement our Bayesian neural networks. MFVI, linearized Laplace and TFSVI were implemented based on the information in the papers, and code for FVI was adapted to the JAX library from the implementation provided by the authors. We further use the GPJAX Python library for experiments involving Gaussian processes (Pinder and Dodd,, 2022).
A.3.8 Hardware

### v3

@@ C.1 Experiments on synthetic data
$y_{i}=\operatorname{sin}(2\pi x_{i})+\epsilon\quad\text{with}\quad\epsilon\sim\mathcal{N}\mathopen{}\left(0,\sigma_{n}^{2}\right)\mathclose{}$ (C.1)
and draw $x_{i}\sim\mathcal{U}([-1,-0.5]\cup[0.5,1])$. When not otherwise specified, we use $\sigma_{n}=0.1$. On the plots, the data points are shown as gray circles, inferred mean functions as red lines, their 2-standard-deviations interval around the mean in light green, and functions sampled from the approximate posterior as green lines. In general, we consider two hidden-layer BNNs with 30 neurons per layer and hyperbolic tangent activation (Tanh) functions. Specifically in figure 11, the small BNN has the same architecture as above while the large BNN has 100 neurons per layer. All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior. The prior scale of TFSVI [Rudner et al., 2022b] is set to $\sigma_{p}=0.2$ and $\sigma_{p}=0.75$ for MFVI [Blundell et al., 2015] and Laplace [Immer et al., 2021]. For the Gaussian process posterior baseline, we fit the prior parameters by maximizing the log-marginal likelihood [Williams and Rasmussen, 2006]. Apart from the cases where the parameters of the GP prior used for GFSVI (our method) and FVI [Sun et al., 2019] are explicitly stated, we consider a constant zero-mean function and find the parameters of the covariance function by maximizing the log-marginal likelihood from mini-batches [Chen et al., 2022]. Except where otherwise stated, we estimate the functional KL divergences with 500 measurement points and use the regularized KL divergence with $\gamma=10^{-10}$.
@@
We sample 100 data points perturbed by Gaussian noise with $\sigma_{n}=0.1$ from the two moons data [Pedregosa et al., 2011]. On the plots, the data points are shown as red (class 0) and blue (class 1) dots. We plot the mean and 2-standard-deviations of the probability that ${\bm{x}}$ belongs to class 1 with respect to the posterior (i.e. $p(y=1\,|\,{\bm{w}}^{(k)},{\bm{x}})$) which we estimate from $K=100$ samples ${\bm{w}}^{(k)}\sim q_{\phi}({\bm{w}})$ for $k=1,\dots,K$. We consider two hidden-layer BNNs with 100 neurons per layer and hyperbolic tangent activation (Tanh) functions. All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior. The prior scale of MFVI [Blundell et al., 2015] is set to $\sigma_{p}=0.8$ and $\sigma_{p}=1.0$ for TFSVI [Rudner et al., 2022b] and Laplace [Immer et al., 2021]. For the Gaussian process posterior baseline, we approximate the intractable posterior using the Laplace approximation and find the prior parameters by maximizing the log-marginal likelihood [Williams and Rasmussen, 2006]. The GP prior for GFSVI (our method) and FVI [Sun et al., 2019] has a constant zero-mean function and we find the parameters of the covariance function by maximizing the log-marginal likelihood from mini-batches [Chen et al., 2022] using the method to transform classifications labels into regression targets from Milios et al. [2018]. We estimate the functional KL divergences with 500 measurement points and use the regularized KL divergence with $\gamma=10^{-10}$.
C.2 Ocean current modeling experiment
Following Cinquin et al. [2024], we apply the Helmholtz decomposition to the neural network $f$ as
$f(\cdot,{\bm{w}})=\operatorname{grad}\Phi(\cdot,{\bm{w}}_{1})+\operatorname{rot}\Psi(\cdot,{\bm{w}}_{2})$ (C.2)
@@ C.2 Ocean current modeling experiment
C.3 Regression experiments with tabular data
@@
We evaluate the predictive performance of our model on regression datasets from the UCI repository [Dua and Graff, 2017] described in table 5. These datasets are also considered in Sun et al. [2019], Wild et al. [2022b] but we include two additional larger ones (Wave and Denmark). We perform 5-fold cross validation, leave out one fold for testing, consider 10% of the remaining 4 folds as validation data and the rest as training data. We report mean and standard-deviation of the average expected log-likelihood and average mean square error on the test fold. We also report the mean rank of the methods across all datasets by assigning rank 1 to the best scoring method as well as any method who’s error bars overlap with the highest score’s error bars, and recursively apply this procedure to the methods not having yet been assigned a rank. The expected log-likelihood is estimated by Monte Carlo integration when it is not available in closed form (MFVI, TFSVI and FVI) with 100 posterior samples. We preprocess the dataset by encoding categorical features as one-hot vectors and standardizing the features and labels.
@@
We compare our GFSVI method to two weight-space inference methods (mean-field variational inference [Blundell et al., 2015] and linearized Laplace [Immer et al., 2021]) and two function-space inference methods (FVI [Sun et al., 2019] and TFSVI [Rudner et al., 2022b]). While FVI uses GP priors, TFSVI performs inference in function space but with the pushforward to function space of the variational distribution and prior on the weights. We compute the function-space (regularized) KL divergence using a set of 500 measurement points sampled from a uniform distribution for GFSVI and TFSVI, and 50 points drawn from a uniform distribution along with 450 samples from the training batch for FVI as specified in Sun et al. [2019]. All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior. We also provide results with a GP [Williams and Rasmussen, 2006] when the size of the dataset allows it, and a sparse GP [Hensman et al., 2013]. As we restrict our comparison to BNNs, we do not consider the GP and sparse GP as baselines but rather as gold-standards. All models have a Gaussian homoskedastic noise model with a learned scale parameter. All the BNNs are fit using the Adam optimizer [Kingma and Ba, 2017] using a mini-batch size of 2000 samples. We also perform early stopping when the validation loss stops decreasing.
@@
Hyper-parameter optimization is conducted using the Bayesian optimization tool provided by Wandb [Biewald, 2020]. BNN parameters are selected to maximize the average validation expected log-likelihood across the 5 cross-validation folds. We optimize over prior parameters (kernel and prior scale), learning-rate and activation function. We select priors for GFSVI, FVI, sparse GP and GP among the RBF, Matérn-1/2, Matérn-3/2, Matérn-5/2, Linear and Rational Quadratic covariance functions. The GP prior parameters used with GFSVI and FVI are selected by maximizing the log-marginal likelihood from batches as proposed by Chen et al. [2022] and done in Sun et al. [2019]. Hyper-parameters for GPs and sparse GPs (kernel parameters and learning-rate) are selected to maximize the mean log-marginal likelihood of the validation data across the 5 cross-validation folds.
@@
C.4 Classification experiments with image data
@@
We further evaluate the predictive performance of our model on classification tasks with the MNIST [LeCun et al., 2010] and Fashion MNIST [Xiao et al., 2017] image data sets. We fit the models on a random subset of 90% of the provided training split, consider the remaining 10% as validation data and evaluate on the provided test split. We repeat this procedure 5 times with different random seeds and report the mean and standard-deviation of the average expected log-likelihood, accuracy and expected calibration error (ECE) of the mean of the predictive distribution on the test set. The expected log-likelihood is estimated by Monte Carlo integration with 100 posterior samples when it is not available in closed form (MFVI, TFSVI and FVI). We estimate the mean of the predictive distribution to compute the accuracy and the ECE with 100 posterior samples. We preprocess the dataset by standardizing the images.
@@
We compare our GFSVI method to the same baselines as for the regression experiments (see C.3). All the BNN baselines have the same architecture and fully-factorized Gaussian approximate posterior. More specifically, we consider a CNN with three convolutional layers (with output channels 16, 32 and 64) before two fully connected layers (with output size 128 and 10). The convolutional layers use $3\times 3$ shaped kernels. Each pair of convolutional layers is interleaved with a max-pooling layer. We consider three different measurement point distributions $\rho$ to estimate the (regularized) KL divergence in GFSVI, FVI and TFSVI: random, random pixel and kmnist. The random measurement point distribution is sampled from by drawing 50% of the samples from the training data batch and 50% of the samples from a uniform distribution over $[p_{min},p_{max}]^{H\times W\times C}$, where $H$, $W$ and $C$ are respectively the height, width and number of channels of the images, and $p_{min}=v_{min}-0.5\times\Delta$ and $p_{max}=v_{max}+0.5\times\Delta$ where $\Delta=v_{max}-v_{min}$ is the difference between the minimal ($v_{min}$) and maximal ($v_{max}$) pixel values of the data set. The random pixel measurement point distribution is taken from Rudner et al. [2022b] and is sampled from by randomly choosing each pixel value among the ones available from the training data batch at the same position in the $28\times 28$ pixel grid. Finally, the kmnist measurement point distribution is also taken from Rudner et al. [2022b] and is drawn from by randomly sampling data points from the Kuzushiji-MNIST (KMNIST) dataset [Clanuwat et al., 2018]. The KMNIST dataset is a collection of 70’000 gray-scale images of size $28\times 28$ which we preprocess by standardizing the images. We sample 25 measurement points when using random, 25 measurement points when using random pixel and 20 when using kmnist. All the BNNs are trained using the Adam optimizer [Kingma and Ba, 2017] using a mini-batch size of 100. We also perform early stopping when the validation loss stops decreasing.
@@
Hyper-parameter optimization is conducted just like for the regression tasks (see C.3). The Gaussian process prior parameters used with GFSVI and FVI are selected by maximizing the log-marginal likelihood from batches [Chen et al., 2022] using the method to transform classifications labels into regression targets from Milios et al. [2018]. We optimize the same hyper-parameters as for the regression experiments with the exception of the additional $\alpha_{\epsilon}$ parameter introduced by Milios et al. [2018] for the function-space VI methods with GP priors (FVI and GFSVI).
C.5 OOD detection
@@
Following the setup from Malinin et al. [2021] we take epistemic uncertainty to be the variance of the mean prediction with respect to samples from the posterior. We consider the test data to be in-distribution (ID) data and a subset of the song dataset [Bertin-Mahieux et al., 2011] of equal length and with an equal number of features as out-of-distribution (OOD) data. We use the same preprocessing as for regression as well as the same baselines with the same hyper-parameters (see section C.3). We first fit a model, then evaluate the extend by which the epistemic uncertainty under the model is predictive of the ID and OOD data using a single threshold obtained by a depth-1 decision tree fit to minimize the classification loss. We report the mean and standard error of the accuracy of the threshold to classify OOD from ID data based on epistemic uncertainty across the 5 folds of cross-validation. We also provide results obtained using a GP and sparse GP as gold standard.
@@
Following the setup by Osawa et al. [2019], we take the epistemic uncertainty to be the entropy of the mean of the predictive distribution with respect to samples from the posterior. We evaluate models trained on MNIST using MNIST’s test split as ID data and a subset of the training set of Fashion MNIST as OOD data. Likewise, we evaluate models trained on Fashion MNIST using Fashion MNIST’s test split as ID data and a subset of the training set of MNIST as OOD data. We use the same preprocessing as for classification, as well as the same baselines with the same hyper-parameters (see section C.4). We first fit a model, then evaluate the extend by which the epistemic uncertainty under the model is predictive of the ID and OOD data using a single threshold obtained by a depth-1 decision tree fit to minimize the classification loss. We estimate mean of the predictive distribution by Monte-Carlo integration using 100 posterior samples. We report the mean and standard error of the accuracy of the threshold to classify OOD from ID data based on epistemic uncertainty for the 5 models trained on different random seeds (see section C.4).
C.6 Variational measure evaluation
We evaluate our inference method by comparing the samples drawn from the exact posterior over functions with the approximate posterior obtained with our method (GFSVI). We follow the setup by Wilson et al. [2022] and we compute the average Wasserstein-2 metric between 1000 samples drawn from a GP posterior with a RBF kernel evaluated at the test points, and samples from the approximate posterior of GFSVI, sparse GP and FVI evaluated at the same points and with the same prior. We consider the Boston, Concrete, Energy, Wine and Yacht datasets for which the exact GP posterior can be computed and use the same preprocessing as for regression (see section C.3). We report the mean and standard error of the average Wasserstein-2 metric across the 5 folds of cross-validation. The Wasserstein-2 metric is computed using the Python Optimal Transport library [Flamary et al., 2021].
@@ C.6 Variational measure evaluation
C.7 Software
We use the JAX [Bradbury et al., 2018] and DM-Haiku [Hennigan et al., 2020] Python libraries to implement our Bayesian neural networks. MFVI, linearized Laplace and TFSVI were implemented based on the information in the papers, and code for FVI was adapted to the JAX library from the implementation provided by the authors. We further use the GPJAX Python library for experiments involving Gaussian processes [Pinder and Dodd, 2022].
C.8 Hardware